# Scaffold-BPE: Enhancing Byte Pair Encoding for Large Language Models with Simple and Effective Scaffold Token Removal

### v2

@@ A.2 Better Uniformity of Learned Embeddings
Prior works have analyzed the embedding space learned by a model (Provilkov et al. 2019) and found that better uniformity prefers a token embedding space that preserves maximal information (Wang and Isola 2020). To demonstrate our Scaffold-BPE can mitigate token frequency distribution imbalance, thus leading to a better-learned token embedding space with better uniformity, we visualize the token embeddings in the 6.7B-parameter models, following Provilkov et al. (2019).
@@ B.1 Model Training Hyper-Parameters
We train three language models with 468M, 1.2B, and 6.7B parameters, respectively. Specifically, the architectures are listed in Table 9. Following LLaMA (Touvron et al. 2023a), we use the AdamW optimizer (Loshchilov and Hutter 2017) with a learning rate of $3.0\times 10^{-4}$, $2k$ warmup steps, and a cosine learning rate decay schedule.
@@ B.2 Evaluation
For fair comparisons, we utilize the open-source pipeline lm-evaluation-harness (Gao et al. 2023) for evaluation. We reported the average evaluation results of the last five checkpoints.

### v3

@@ A.2 Better Uniformity of Learned Embeddings
Prior works have analyzed the embedding space learned by a model Provilkov et al. (2019) and found that better uniformity prefers a token embedding space that preserves maximal information Wang and Isola (2020). To demonstrate our Scaffold-BPE can mitigate token frequency distribution imbalance, thus leading to a better-learned token embedding space with better uniformity, we visualize the token embeddings in the 6.7B-parameter models, following Provilkov et al. (2019).
@@ B.1 Model Training Hyper-Parameters
We train three language models with 468M, 1.2B, and 6.7B parameters, respectively. Specifically, the architectures are listed in Table 9. Following LLaMA (Touvron et al., 2023a), we use the AdamW optimizer (Loshchilov and Hutter, 2017) with a learning rate of $3.0\times 10^{-4}$, $2k$ warmup steps, and a cosine learning rate decay schedule.
@@ B.2 Evaluation
For fair comparisons, we utilize the open-source pipeline lm-evaluation-harness (Gao et al., 2023) for evaluation. We reported the average evaluation results of the last five checkpoints.